If EBIT is $1.3 million, net income is $750,000, and sales are $18 million, what is the operating margin?

A.

4.17%

B.

9.32%

C.

12.56%

D.

18.12%

A bond maturing in 10 years at a par value of $1,000 has a coupon rate of 5.5% and current yield of 7.5%. What is the price of the bond?

A.

$861.04

B.

$924.78

C.

$982.45

D.

$1,015.68

The company balance sheet gives a "snapshot" of a company's assets and who has claims on them on a given date.

True

False

Trend Analysis refers to an analysis of a firm's financial ratios over time.

True

False
Answer #1

1)

Operating margin =Net income/Net sales =0.75/18 = 4.17%

2)

=PV(7.5%,10,5.5%*1000,1000)

=861.04

3)

False, balance sheet does not provide enough data to ascertain the claims of people on assets

4)

True.